OpenStudy (anonymous):

Find delta such that |square root of (5x +1)-4>0.5 if 0<|x-3|

OpenStudy (anteater):

Please double check. Is: \[\left| \sqrt{5x+1}-4 \right| > 0.5\] ?

OpenStudy (anteater):

Or is \[\left| \sqrt{5x+1}-4 \right|<0.5\] ?